OpenStudy (dmezzullo):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3. y − 6 = −3x − 3 y − 6 = −3(x + 1) y = −3x + 3 3x + y = 3

OpenStudy (anonymous):

The equation of a line with slope \(m\) that goes through the point \((x_0,y_0)\) is given by: \[ y-y_0 = m (x-x_0) \]

OpenStudy (anonymous):

the line that passes through the point (−1, 6) and has a slope of −3. \((x_0,y_0)=(-1,6)\) and \(m=3\)

OpenStudy (dmezzullo):

would it be possible to put it in understandable terms?

OpenStudy (aripotta):

mezzfezz, do you know what point-slope form is?

OpenStudy (dmezzullo):

Errrm

OpenStudy (anonymous):

Replace \(m\) with \(3\) then replace \(x_0\) with \(-1\), etc...

OpenStudy (anonymous):

Whoops, \(m\) is replaced with \(-3\)

OpenStudy (dmezzullo):

y – y1 = m(x – x1)

OpenStudy (aripotta):

so x1 and y1 will be replaced with the x and y of the given point (-1, 6)

OpenStudy (dmezzullo):

yes

OpenStudy (aripotta):

and m is the slope

OpenStudy (dmezzullo):

So B.

OpenStudy (aripotta):

so do you know what the answer is now

OpenStudy (dmezzullo):

BBBBBBBBBBBB

OpenStudy (dmezzullo):

RIGHT

OpenStudy (aripotta):

yes :]

OpenStudy (dmezzullo):

AYE MEZZFEZZ IS RIGHT!

OpenStudy (dmezzullo):

ty im done now